OpenStudy (anonymous):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost.Which equation represents the situation? A) y-2=-10(x-25) B) y-4=10(x-45) C) y-65=-10(x-6) D) y-85=10(x-8)

OpenStudy (anonymous):

@austyfoxy

OpenStudy (anonymous):

The answer is B

OpenStudy (anonymous):

Can ya help with one more please?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Thank you!! Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. x=Time babysitting and y= amount earned (4,33) and (8,61) Which equation represents the situation? A) y=1/7x-5 B) y=1/7x+5 C) y=7x-5 D) y=7x+5

OpenStudy (anonymous):

Yeah im pretty sure its D

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Yeah man. Anytime